OpenStudy (anonymous):

u are given 88 feet of fence what woulkd the formula of the area be in terms of lenght of one of its sides?

OpenStudy (anonymous):

any luck ?

OpenStudy (anonymous):

?

OpenStudy (radar):

2a+2b=88 or a+b+44 then a=44-b area is 44b-b^2 Assuming a rectangle with sides a and b